# A 44-minute periodic radio transient in a supernova remnant

### v1

@@ Abstract
Long-period radio transients (LPTs) are a newly discovered class of radio emitters with yet incomprehensibly long rotation periods, ranging from minutes to hours. The astrophysical nature of their isolated counterparts remains undetermined. We report a new LPT, DART J1832-0911 (2656.23±0.15 s period), the first evidence associating such objects to supernova remnants (SNRs). Its dispersion measure distance aligns well with the distance of the SNR, confirming its origin from a supernova explosion. The source displays either phase-locked circularly polarized emission or nearly 100% linear polarization in radio bands. No detectable optical counterpart was found, even with a 10m-class telescope. The J1832-0911’s SNR association, stable, highly polarized emission, and abnormally long period strongly favor its origin from a young neutron star, whose spin has been braked, possibly by interaction with supernova’s fallback materials. This discovery provides critical insights into the nature of ultra-long period transients and their evolutionary link to stellar remnants.
@@
Long-period radio transients (LPTs) are a new class of periodic radio emitters, characterized by ultra-long rotation periods and strong magnetic fields. These features place LPTs significantly outside the conventional framework of rotation-powered neutron star pulsars or white dwarfs (?, ?, ?, ?). As for their isolated stellar candidates, the two primary models—spin-down powered pulsars and white dwarfs—each face significant challenges in explaining the observed emissions if they independently evolve in a traditional way: losing rotation energy through radio emission. Neutron stars with ultra-long periods typically fall below the so-called “death line”, where their rotational energy becomes insufficient to sustain coherent radio emission (?, ?, ?), while white dwarfs often have insufficient magnetic fields to produce the observed periodic, coherent emissions (?). Among reported LPT cases, four are identified as dwarf binary systems where accretion processes plays a role in driving the observed emissions (?, ?, ?, ?), while four are isolated sources: GLEAM-X J162759.5-523504.3 (18.18 min) (?), GPM J1839-10 (21 min) (?), ASKAP J1935+2148 (53 min) (?), CHIME J0630+25 (7 min) (?). Although theoretical models lean toward isolated LPTs being magnetars (i.e. neutron stars with strong magnetic field), direct observational evidence and evolution studies remain limited.
DART Observation
@@
We report a new LPT in a supernova remnant, DART J1832-0911, detected by DAocheng Radio Telescope (DART) (see supplementary material) with interferometric imaging across a frequency range of 149-459 MHz. The first pulse was detected on Modified Julian Date (MJD) 60374. From MJD 60401 and MJD 60443, we conducted continuous observations of the transient using an interferometric imaging technique, with full polarization and at double center frequency. The recorded pulses showed an estimated peak flux density between 0.5–2 Jy. After this period, the transient entered a long-period quiescent state, as recorded by DART observation.
@@
The interferometric image allowed us to determine the celestial coordinates to be RA=18h32m48.8s () and DEC=-09d11m34s (), situating it within the projected region of supernova remnant G22.7-0.2. Top panel in Fig.1 displays the pulse-on image of DART J1832-0911 at 425 MHz. The bottom panel shows the observed single pulses in two Stokes components. Additional pulse profile shapes are provided in the supplementary material (Fig.S1). All pulse series presented in this study have been aligned through periodic folding and corrected for delay time. By analyzing the time of arrival (TOA) of single pulses recorded by DART, we derive a spin period of s with an estimated period derivative of s/s. Fig.2 displays the - diagram of DART J1832-0911 as well as other LPTs.
FAST Observation
@@
To further investigate, we used the Five-hundred-meter Aperture Spherical Telescope (FAST) to observe the source at L-band during DART’s non-detection period on MJD 60559. During a 1-hour observation, we detected a short pulse with a peak flux density of approximately 94 mJy. The duration of the pulse (i.e. the pulse width) is 0.2 s, just 0.007% of the rotation period. This could point to a highly energetic and sharp emission event occurring at the strong magnetic field, like the dwarf pulse emission process of a neutron star (?, ?).
@@
Optical and X-ray Counterpart Search
We conducted an optical observation using the Gran Telescopio Canarias (GTC) on 30 July , 2024. The integration time for this observation was 1.25 hours. We conformed no counterparts in optical observation.
An r-band image of the DART source location was obtained using the broad-band imaging mode of OSIRIS on the GTC. As shown in Fig.4, there are two optical candidates within the designated zone. However, the light curves of these candidates show no periodic variations on timescales of tens of minutes, indicating they are not associated with DART J1832-0911. Given an upper limit of 26.6 magnitudes for the GTC observations and assuming a white dwarf radius of at a distance of 4.5 kpc, any WD detectable by the GTC would have an effective temperature below approximately , indicating an evolution age at hundreds of millions years (see supplementary material). The absence of an optical counterpart supports the possibility of a compact neutron star rather than a WD. Additionally, our analysis shows that a spin-down powered WD would be unable to account for the observed emission power of a transient with characters like those of DART J1832-0911 (see supplementary material).
We searched for an X-ray counterpart using archival data from the XMM-Newton observatory, obtained on March 13, 2011 (Observation ID: 0654480101). The observation had an exposure time of 1.54 × 104 s and covered the source with both XMM-PN and XMM-MOS2 detectors in the 0.2-10keV. Within the positional uncertainty of DART J1832-0911, only a single photon was detected by XMM-PN (see Fig.S8 in supplementary materia). Consequently, no X-ray counterpart was identified for either the transient or the associated SNR region. This yields an upper limit of 6.648x10-14 ergs cm2 s-1 in 0.2-10 keV at a 99% confidence level.
We suggest that the absence of X-ray detection may result from an unsuitable temporal observation window for the history data or inherently weak or absent X-ray emission from this source.
Dispersion Measure and SNR Association Analysis
We calculated the dispersion measure (DM) using dual-frequency pulse series from DART and the short pulse spectra of FAST. By relating the TOA delay to the dispersion effect observed at two simultaneous frequencies, we obtained a DM of pc . A DM search of the FAST spectra yielded pc . These measurements from the two observatories are consistent with the 1 sigma error. Given the broader bandwidth and finer temporal resolution of FAST, we consider this result to be the precise DM for DART J1832-0911.
Based on this DM, we estimated the distance to DART J1832-0911, finding it consistent with the distance of SNR G22.7-0.2. Previous studies estimate the distance to SNR G22.7-0.2 at 4.4 kpc (?) or 4.7 kpc (?), based on molecular cloud analyses. Meanwhile, the DM-derived distance to DART J1832-0911, using the YMW16 electron-density model (?), is approximately 4.5 kpc, suggesting that this LPT resides within the SNR bubble. No other compact objects have been detected within this SNR.
We also investigated the X-ray emission of SNR G22.7-0.2 using data from XMM-Newton and previous studies. Our analysis, consistent with earlier research, reveals no evidence of extended X-ray emission within the radio range. The cessation of X-ray emission from the SNR suggests that the remnant’s shock velocity, , has decreased below 200 km/s—insufficient to heat the gas to X-ray-emitting temperatures (?, ?). Consequently, we can reasonably infer that the SNR has entered the radioactive phase, during which the age of the SNR can be estimated years where pc is the SNR radius. Assuming a lower limit for the shock velocity of 30 km/s (?), we obtain an upper age limit of years.
Radio Emission Properties
DART J1832-0911 displays a range of emission characteristics during its active radio period. Prior to either undetectable weak emission or nulling, the source undergoes mode changes modulated by variations in pulse width and strength, as illustrated in the bottom panel of Fig.1. These time-dependent mode changes reveal an evolution from bright, wide pulses to weaker, narrower ones. In the wide mode, pulse widths range from 100 s to 300 s, with a flux density of about 2 Jy. In the narrow mode, pulse width narrows to approximately 40 s, and flux density increases to roughly 1 Jy. During the emission-quiescent state, as observed by DART, we captured a short-lived pulse by FAST. Within the expected periodic phase, this pulse lasted 0.2 s and exhibited a peak flux of mJy. Such brief, isolated emission components have also been observed in other long-period transients (?, ?).
Stellar Origin and Evolution of DART J1832-0911
The discovery of DART J1832-0911 strongly favors a young neutron star as one origin of long period radio transients. The spatial association between this source and the SNR indicates that it is likely the stellar remains of a supernova explosion, specifically a neutron star rather than a white dwarf. Given the characteristic age of both DART J1832-0911 and SNR G22.7-0.2, it is reasonable to conclude that DART J1832-0911 is a young neutron star. The emission properties further suggest a pulsar-like emitter of DART J1832-0911. The narrow, phase-locked CP indicates a stable geometric structure of the magnetic field of DART J1832-0911 even at small scales. This stability argues against models that rely on dynamic magnetic fields as the primary driver of radio emission, such as star flares (?) and dwarf binary systems (?, ?). The nearly 100% linearly polarized burst signals support both theoretical radio emission mechanism and observation related to magnetars (?, ?, ?, ?), which is believed to be a kind of young neutron stars (?). Such short-lived pulse emitted during a emission-dull of -quiescent state is similar to the dwarf pulse observed during pulsars’ nulling period (?, ?, ?, ?, ?). These emissions are believed to arise from rain-like drops of particles formed via pair production near the neutron star surface (?, ?).
A neutron star nature of LPTs significantly challenges conventional evolution models of fast spinning neutron star scenarios (?), which typically have rotation periods of less than a few minutes, as illustrated in Fig.2. Theoretically, neutron stars—formed as remnants of supernova explosions—are born with fast spin periods on the order of milliseconds if they inherit most or all of the angular momentum from their progenitors. Consequently, the emergence of LPTs suggests that these stars could form under two possible conditions: either with a significant loss of angular momentum from their progenitors such as neutrino emission (?, ?), or through some rotational braking process during their spin evolution (see (?) for a review).
Based on the fact that DART J1832-0911 is surrendered by supernova remnant matters, the braking process of a neutron star spun down by its surrounding supernova fallback disc (?, ?, ?) should be the most reasonable explanation for the long period. We simulate the reasonable period range based on a fallback disc model (?), as shown in Fig.2. The results suggest that the observed LPTs could be well explained in the period between and years after the supernova explosion. Therefore, the association between DART J1832-0911 and SNR G22.7-0.2, with an age estimate ranging from to years, could further support the hypothesis that such LPTs may originate from interactions between young neutron stars and fallback discs.
@@
| eak flux density (Jy) | 0.5-2 | 0.0960.027 |
Supplementary materials
Materials and Methods
Figs. S1 to S8
Tables S1
References (7-0)
Supplementary Materials for

### v2

@@ Abstract
[…]
Keyword: radio astronomy; long period radio transient; supernova remnant; radio pulsar
These authors contributed equally to this work
* Corresponding authors: yanjingye@nssc.ac.cn
I Introduction
@@
Long-period radio transients (LPTs) are a new class of periodic radio emitters, characterized by ultra-long rotation periods and strong magnetic fields. These features place LPTs significantly outside the conventional framework of rotation-powered neutron star pulsars or white dwarfs (marsh2016radio; pelisoli20235; hurley2022radio; hurley2023long; caleb2024emission; rea2024long; de2025sporadic; lee2025emission; dong2025chime; bloot2025strongly). Among reported LPT cases, several are identified, or argued to originate from, white-dwarf binary systems marsh2016radio; pelisoli20235; de2025sporadic; hurley20242; dong2025chime; bloot2025strongly, including both spin-down powered white-dwarf pulsars (i.e. AR Sco) and accretion-powered systems. Another group shows no compelling evidence for companions: GLEAM-X J162759.5-523504.3 (18.18 min) hurley2022radio, GPM J1839-10 (21 min) hurley2023long, ASKAP J1935+2148 (53 min) caleb2024emission, CHIME J0630+25 (7 min) dong2024discovery. For the stellar candidates LPTs without confirmed companions, the two primary models—spin-down powered pulsars and white dwarfs—each face significant challenges in explaining the observed emissions if they independently evolve in a traditional way: losing rotation energy through radio emission. Neutron stars with ultra-long periods typically fall below the so-called “death line”, where their rotational energy becomes insufficient to sustain coherent radio emission chen1993pulsar; young1999radio; caleb2022discovery, while white dwarfs often have insufficient magnetic fields to produce the observed periodic, coherent emissions rea2024long. Although theoretical models lean toward isolated LPTs being magnetars (i.e. neutron stars with strong magnetic field) cooper2024beyond; men2025highly; beniamini2023evidence, direct observational evidence and evolution studies remain limited.
II Observation and Results
II.1 DART Observation
@@
We report a new LPT in a supernova remnant, DART J1832-0911, detected by DAocheng Radio Telescope (DART) (see supplementary material) with interferometric imaging across a frequency range of 149-459 MHz. The first pulse was detected on Modified Julian Date (MJD) 60374. From MJD 60401 to MJD 60443, we conducted continuous observations of the transient using an interferometric imaging technique with full polarization, simultaneously recording data at two distinct center frequencies, each with a bandwidth of 4 MHz. The recorded pulses showed an estimated peak flux density between 0.5–2 Jy. After this period, the transient entered a long-period quiescent state, as recorded by DART observation.
@@
The interferometric image allowed us to determine the celestial coordinates to be RA=18h32m48.5s () and DEC=-09d11m17s (), situating it within the projected region of supernova remnant G22.7-0.2. Top panel in Fig.1 displays the pulse-on image of DART J1832-0911 at 425 MHz. The bottom panel shows the observed single pulses in two Stokes components. Additional pulse profile shapes are provided in the supplementary material (Fig.S1). All pulse series presented in this study have been aligned through periodic folding and corrected for delay time. By analyzing the time of arrival (TOA) of single pulses recorded by DART, we derive a spin period of s with an estimated period derivative of s/s. Fig.2 displays the - diagram of DART J1832-0911 as well as other LPTs.
II.2 FAST Observation
@@
To further investigate, we used the Five-hundred-meter Aperture Spherical Telescope (FAST) to observe the source at L-band during DART’s non-detection period on MJD 60559. During a 1-hour observation, we detected a short pulse with a peak flux density of approximately 94 mJy. The duration of the pulse (i.e. the pulse width) is 0.2 s, just 0.007% of the rotation period. This could point to a highly energetic and sharp emission event occurring at the strong magnetic field, like the dwarf pulse emission process of a neutron star chen2023strong; yan2024dwarf.
@@
| Peak flux density (Jy) | 0.5-2 | 0.0960.027 |
II.3 Optical and X-ray Counterpart Search
We conducted an optical observation using the Gran Telescopio Canarias (GTC) on 30 July 2024. The integration time for this observation was 1.25 hours.
No credible optical counterpart was identified in the GTC observation, effectively ruling out any pulsating white dwarf with a temperature 17,000K as its origin. An r-band image of the DART source location was obtained using the broad-band imaging mode of OSIRIS on the GTC. As shown in Fig.4, there are four optical candidates within the designated zone. However, the light curves of these candidates show no periodic variations on timescales of tens of minutes, indicating they are not associated with DART J1832-0911. Given an upper limit of 26.6 magnitudes for the GTC observations and assuming a white dwarf radius of at a distance of 4.5 kpc, any WD detectable by the GTC would have an effective temperature below approximately , indicating an evolution age of hundreds of millions years (see supplementary material). The absence of an optical counterpart supports the possibility of a compact neutron star rather than a WD. Additionally, our analysis shows that a spin-down powered WD would be unable to account for the observed emission power of a transient with characteristics like those of DART J1832-0911 (see supplementary material).
We searched for any possible X-ray counterpart using archival data from the XMM-Newton observatory, obtained on March 13, 2011 (Observation ID: 0654480101). The observation had an exposure time of s and covered the source with both XMM-PN and XMM-MOS2 detectors in the 0.2-10 keV. Within the positional uncertainty of DART J1832-0911, only a single photon was detected by XMM-PN. Consequently, no X-ray counterpart was identified for either the transient or the associated SNR region according to the observation in 2011. This yields an upper limit of erg cm-2 s-1 in 0.2-10 keV at a 99% confidence level.
Independently of our work, a recent study wang2025detection has reported a 44-min periodic radio transient in the same direction, consistent with our detection (referred to as ASKAP J18320911 in their work). The authors performed a comprehensive X-ray analysis of J18320911 using additional instruments, including Chandra, Swift, and Einstein Probe Yuan2015EP, leading to the identification of a transient X-ray counterpart detected by Chandra. Their observations revealed an extreme contrast between a bright X-ray outburst ( erg s-1) and a deep quiescent state ( erg s-1). Assuming a distance of 4.5 kpc, these correspond to flux limits of approximately erg cm-2 s-1 in the 1–10 keV band. We suggest that rapid follow-up X-ray observations conducted during the radio-bright phases of future detected LPTs will be critical for unveiling their physical nature.
Archival XMM-Newton observations from 2011 revealed no detectable X-ray emission from DART J1832–0911, whereas an X-ray burst was detected by Chandra in February 2024 but not in August (wang2025detection), indicating that J1832–0911 is an intermittent X-ray transient.
II.4 Dispersion Measure and SNR Association Analysis
We calculated the dispersion measure (DM) using dual-frequency pulse series from DART and the short pulse spectra of FAST. By relating the TOA delay to the dispersion effect observed at two frequencies simultaneously recorded, we obtained a DM of pc . A DM search of the FAST spectra yielded . These measurements from the two observatories are consistent within the 3 sigma error. Given the broader bandwidth and finer temporal resolution of FAST, we consider this result to be the precise DM for DART J1832-0911.
We estimated the DM distance of DART J1832-0911, it is consistent with the distance of SNR G22.7-0.2. Previous studies estimate the distance to SNR G22.7-0.2 at 4.4 kpc su2014interaction or 4.7 kpcranasinghe2018revised, based on molecular cloud analyses. Meanwhile, the DM-derived distance to DART J1832-0911 was estimated using three Galactic free-electron density models: YMW16 yao2017new, NE2001 cordes2002ne2001, and NE2025 ocker2026ne2025, yielding distances of approximately 4.5 kpc, 5.9 kpc, and 6.4 kpc, respectively. All three distances are broadly consistent with the SNR distance given the known systematic uncertainties of free-electron density models, which can reach 30–50%. This suggests that J1832-0911 may reside within the SNR bubble. Nevertheless, the DM-estimated distance remains debated due to the electron-density model dependence.
To evaluate the reliability of DM-based distance estimation along this line of sight, we compare DM-derived distances with VLBI parallax distances with two pulsars deller2019microarcsecond,which are located within in declination of DART J18320911: PSR J18200427 and PSR J19010906. For PSR J18200427 (DM pc cm-3), the NE2001, YMW16, and NE2025 models yield distances of 1.94, 2.92, and 2.01 kpc, respectively, while a VLBI parallax distance is kpc. For PSR J19010906 (DM pc cm-3), the DM models give 2.13, 2.89, and 2.43 kpc, and the VLBI distance is kpc. We estimate the uncertainty in DM-derived distances based on the VLBI measurements using two approaches. First, we evaluate each of the three DM models individually. The average fractional deviations between the DM distances and the VLBI distances for the two sightlines are 20% for NE2001, 25% for YMW16, and 27% for NE2025. Based on these model-specific uncertainties, the DM-derived distances of DART J18320911 are 4.7–7.1 kpc (NE2001), 3.4–5.6 kpc (YMW16), and 4.7–8.1 kpc (NE2025). Taking the full range of these estimates, we adopt a DM distance of 3.4 to 8.1 kpc. The second estimation method combines all three DM models. We take the mean of the three model predictions as the reference distance, kpc. The total uncertainty is estimated by combining in quadrature the mean single-model systematic uncertainty (24%) and the fractional model-to-model dispersion (), yielding a combined fractional uncertainty of , or 2.4 kpc. This gives a consolidated distance range for DART J18320911 of 3.2–8.0 kpc. The independently determined distance to SNR G22.70.2 of 4.4–4.7 kpc su2014interaction; ranasinghe2018revised falls within both the individual-model ranges and the consolidated interval, supporting a physical association between DART J18320911 and the SNR.
[…]
We inspected the CO observations toward SNR G22.7-0.2. The velocity of CO emission ranges from 75 to 79 km/s su2014interaction, consistent with the absorption ( km/s) reported in Wang et al. wang2025detection. Such alignment in the gas dynamics strongly supports that the compact object and the SNR reside within the same large-scale ISM structure.
We inspect all compact objects within the 16 deg2 region FoV of our DART observation. There are 17 other sources, which fall within the DM distance of 2 kpc to 8 kpc and an age below 1 million years, accommodating a 100% uncertainty allowed for the DM uncertainty and the SNR distance, and a 1000% uncertainty allowed for the upper limits of the SNR age, as indicated in the Fig.5. The random chance of any of these sources to get within the distance between the LPT and the SNR center is lower than 0.7% (see supplementary material). Note that these sources have a chance of being SNR related. The contamination from the field pulsar population has to be much lower. And the previous pulsar (PSR) -SNR association studies reported no case of pulsars discovered around SNR to be later revealed to a chance coincidence (see supplementary material for a detail discussion).
We also investigated the X-ray emission of SNR G22.7-0.2 using data from XMM-Newton and previous studies. Our analysis, consistent with earlier research, reveals no evidence of extended X-ray emission within the radio range . The cessation of X-ray emission from the SNR suggests that the remnant’s shock velocity, , has decreased below 200 km/s—insufficient to heat the gas to X-ray-emitting temperatures mckee1974x; chevalier1999supernova. Consequently, we can reasonably infer that the SNR has entered the radiative phase, during which the age of the SNR can be estimated as years where pc is the SNR radius. Assuming a lower limit for the shock velocity of 30 km/s which is the minimum shock velocity required to produce the observed high-velocity wings and redshifted broadening in the CO emission of this SNR su2014interaction, we obtain an upper age limit of years. Assuming that DART J18320911 is associated with SNR G22.70.2, we use the age and radius of the SNR to estimate its projected transverse velocity to be 26–176 km/s (see Section S6 for details), which aligns well with the velocity range of the SNR-associated pulsarsmalov2021pulsars.
Conclusively, the widely used association judgement: positional overlap, distance, proper motion, and the consistency between the SNR age and the spin-down timescale expected from fallback disk models, shows no evidence against the physical association between DART J1832-0911 and SNR G22.7-0.2. Nevertheless, further strong evidence from direct observational measurements is required .
II.5 Radio Emission Properties
DART J1832-0911 displays a range of emission characteristics during its active radio period. Prior to either undetectable weak emission or nulling, the source undergoes mode changes modulated by variations in pulse width and strength, as illustrated in the bottom panel of Fig.1. These time-dependent mode changes reveal an evolution from bright, wide pulses to weaker, narrower ones. In the wide mode, pulse widths range from 100 s to 300 s, with a flux density of about 2 Jy. In the narrow mode, pulse width narrows to approximately 40 s, and flux density decreases to roughly 0.5 Jy. During the emission-quiescent state, as observed by DART, we captured a short-lived pulse by FAST. Within the expected periodic phase, this pulse lasted 0.2 s and exhibited a peak flux of mJy. Such brief, isolated emission components have also been observed in other long-period transients hurley2023long; caleb2024emission.
III Conclusion
We report the discovery of a long-period radio transient, DART J18320911, detected with interferometric imaging by the Daocheng Radio Telescope (DART) across 149–459 MHz. The spin period of this LPT is s (44.27 min). The interferometric localization (RA , Dec , ) situates DART J18320911 within the projected extent of the supernova remnant G22.70.2. During the DART non-detection epoch, a short pulse detected by FAST at L-band occurred at the predicted rotational phase and is nearly linearly polarized with a well-measured RM, demonstrating that the source can emit extremely narrow, highly polarized bursts despite long intervals of radio quiescence.
The DM measurements of J1832-0911 from DART and FAST are roughly consistent, and the DM-derived distance (, and kpc from YMW16, NE2001, and NE2025 models, respectively) are broadly consistent with independent distance estimates for SNR G22.70.2, supporting a physical association between the compact object and the remnant environment. Complementary CO kinematics and HI absorption arguments further strengthen the case that both the SNR and the transient reside in the same large-scale ISM structure. A field-population analysis yields a low probability for a chance alignment with unrelated compact objects, adding statistical weight to the association scenario.
DART J18320911 shows distinctive radio emission and polarization phenomenology. In DART bands, the activity evolves between wider and narrower emission modes and ultimately into a long quiescent state, with peak flux densities of order 0.5–2 Jy during active epochs. Polarization properties are notably intermittent: many pulses are weakly polarized, while (i) narrow, phase-locked circular polarization components appear in the wide mode with reaching and occasional sign reversals, indicating a stable and geometrically confined emission region; and (ii) the FAST burst shows nearly pure linear polarization with a flat position angle across the burst window, pointing to an ordered magneto-ionic geometry during that event.
Multi-wavelength follow-up and archival searches impose additional constraints. Optical observation reveals no convincing counterpart down to a deep limiting magnitude, disfavoring a luminous white-dwarf binary interpretation. Archival XMM–Newton observations from 2011 yielded no detectable X-ray counterpart, setting a 99% upper limit on the 0.2–10 keV flux. Together with the reported Chandra detection of X-ray bursts and subsequent non-detection, this suggests that the source can enter a deep X-ray quiescent state and motivates rapid, triggered X-ray follow-up during future radio-bright windows.
IV DISCUSSION
DART J1832-0911 provides several important advances over previously reported LPTs. First, it is the strongest current case for a physical association between an LPT and a supernova remnant, thereby providing the clearest observational support so far that at least some isolated LPTs originate from young neutron stars. Second, the plausible SNR association supplies an external age constraint, allowing the ultra-long spin period to be discussed in an evolutionary context and making post-supernova braking scenarios, such as fallback-disk interaction, substantially more plausible. Third, the source exhibits distinctive polarization behaviour, including phase-locked circular polarization and a high linearly polarized dwarf pulse, which places new constraints on the radio-emission geometry and further distinguishes this source from LPTs associated with white-dwarf binary or accretion-powered systems. Taken together, these properties favor a young neutron star as one origin of at least a subset of long-period radio transients.
The spatial association between this source and the SNR indicates that it is likely the stellar remains of a supernova explosion, specifically a neutron star rather than a white dwarf. Given the characteristic age of both DART J1832-0911 and SNR G22.7-0.2, it is reasonable to suggest that DART J1832-0911 might be a young neutron star. The emission properties further suggest a pulsar-like emitter of DART J1832-0911. The narrow, phase-locked CP indicates a stable geometric structure of the magnetic field of DART J1832-0911 even at small scales. This stability argues against models that rely on dynamic magnetic fields as the primary driver of radio emission, such as star flares benz2010physical and dwarf binary systems marsh2016radio; de2025sporadic. The nearly 100% linearly polarized burst signals support both theoretical radio emission mechanism and observation related to magnetarscamilo2006transient; camilo2007polarized; kramer2007polarized; levin2012radio, which is believed to be a class of young neutron stars kaspi2017magnetars. Such short-lived pulse emitted during an emission-dull or -quiescent state is similar to the dwarf pulse observed during pulsars’ nulling periodchen2023strong; yan2024dwarf; srostlik2005core; kloumann2010long; esamdin2012psr. These emissions are believed to arise from rain-like drops of particles formed via pair production near the neutron star surfacechen2023strong; yan2024dwarf. It is worth noting, however, that in the case of LPTs, such emissions may occur regularly during both emission-loud and -quiet phases, but remain undetected due to the sensitivity and temporal resolution limitations of DART.
A neutron star assumption of LPTs significantly challenges conventional evolution models of fast spinning neutron star scenariosrea2024long, which typically have rotation periods of less than a few minutes, as illustrated in Fig.2. Theoretically, neutron stars—formed as remnants of supernova explosions—are born with fast spin periods on the order of milliseconds if they inherit most or all of the angular momentum from their progenitors. Consequently, the emergence of LPTs suggests that these stars could form under two possible conditions: either with a significant loss of angular momentum from their progenitors such as neutrino emission epstein1978neutrino; camelio2016spin, or through some rotational braking process during their spin evolution (see heger2005presupernova for a review).
Given that DART J18320911 is embedded in the SNR environment, rotational braking by a supernova fallback disk menou2001stability; ronchi2022long; yang2024instability provides the most natural explanation for its ultra-long spin period. We simulate the reasonable period range based on a fallback disc model yang2024instability, as shown in Fig.2. The results suggest that the observed LPTs could be well explained in the period between and years after the supernova explosion. Therefore, the association between DART J1832-0911 and SNR G22.7-0.2, with an age estimate ranging from to years, could further support the hypothesis that such LPTs may originate from interactions between young neutron stars and fallback discs.
Acknowledgements.
This work is supported by the Strategic Priority Research Program of the Chinese Academy of Sciences.Supplementary data for